A train travels 150 kilometers in the same time it takes a plane to travel 400 Kilometers. If the speed of the plane is 20 kilometers per hour less than 3 times the speed of the train , find both speeds?
Answer by checkley71(8403)

150/x=400/3(x-20) or 400x=450x-9000 or 50x=9000 or x=180 speed of the train &
3(180-20) or 3*160=480 speed of the plane